# biblio.bib

### v1

@@ 4.3 Uniform fluctuation estimates
$\delta^{h,i}_{t}=\int_{0}^{t}[b_{s_{h}}(X^{h,i},\pi(\vec{X}^{h}))-b_{s_{h}}(\overline{X}^{h,i},\L(\overline{X}^{h,i}))]\mathrm{d}s\\ +\int_{0}^{t}[\sigma_{s_{h}}(X^{h,i},\pi(\vec{X}^{h}))-\sigma_{s_{h}}(\overline{X}^{h,i},\L(\overline{X}^{h,i}))]\mathrm{d}B^{i}_{s}.$
@@
$\mathbb{E}\big{|}b_{s_{h}}(X^{h,i},\pi(\vec{X}^{h}))-b_{s_{h}}(\overline{X}^{h,i},\L(\overline{X}^{h,i}))\big{|}^{2}\leq C\mathbb{E}\big{[}\sup_{0\leq r\leq s_{h}}|X^{h,i}_{r}-\overline{X}^{h,i}_{r}|^{2}\big{]}+CN^{-1}.$
@@
$\mathbb{E}\sup_{0\leq s\leq t}\bigg{|}\int_{0}^{s}[\sigma_{r_{h}}(X^{h,i},\pi(\vec{X}^{h}))-\sigma_{r_{h}}(\overline{X}^{h,i},\L(\overline{X}^{h,i}))]\mathrm{d}B^{i}_{r}\bigg{|}^{2}\\ \leq C\int_{0}^{t}|\sigma_{s_{h}}(X^{h,i},\pi(\vec{X}^{h}))-\sigma_{s_{h}}(\overline{X}^{h,i},\L(\overline{X}^{h,i}))|^{2}\mathrm{d}s,$
@@
$b_{s}^{h,i}:=b_{s_{h}}(X^{h,i},\pi(\vec{X}^{h}))-b_{s_{h}}(\overline{X}^{h,i},\L(\overline{X}^{h,i}))-[b_{s}(X^{0,i},\pi(\vec{X}^{\infty}))-b_{s}(\overline{X}^{\infty,i},\L(\overline{X}^{0,i}))],$
$\sigma_{s}^{h,i}:=\sigma_{s_{h}}(X^{h,i},\pi(\vec{X}^{h}))-\sigma_{s_{h}}(\overline{X}^{h,i},\L(\overline{X}^{h,i}))-[\sigma_{s}(X^{0,i},\pi(\vec{X}^{\infty}))-\sigma_{s}(\overline{X}^{\infty,i},\L(\overline{X}^{0,i}))].$
@@
$+b_{s_{h}}(\overline{X}^{h,i},\pi(\vec{\overline{X}}^{h}))-b_{s_{h}}(\overline{X}^{h,i},\L(\overline{X}^{h,i}))-\big{[}b_{s}(\overline{X}^{h,i},\pi(\vec{\overline{X}}^{h}))-b_{s}(\overline{X}^{h,i},\L(\overline{X}^{h,i}))\big{]}$
@@
$+b_{s}(\overline{X}^{h,i},\pi(\vec{\overline{X}}^{h}))-b_{s}(\overline{X}^{h,i},\L(\overline{X}^{h,i}))-\big{[}b_{s}(\overline{X}^{i,0}_{s},\pi(\vec{\overline{X}}^{0}_{s}))-b_{s}(\overline{X}^{i,0},\L(\overline{X}^{i,0}))\big{]}.$
@@
$\max_{0\leq r\leq 1}\mathbb{E}\big{[}\big{|}\partial_{s}b_{(1-r)s_{h}+rs}(\overline{X}^{h,i},\pi(\vec{\overline{X}}^{h}))-\partial_{s}b_{(1-r)s_{h}+rs}(\overline{X}^{h,i},\L(\overline{X}^{h,i}))\big{|}\big{]}\leq CN^{-1/2}.$
@@
$\mathrm{d}\delta X^{h}_{t_{h}}$ $=\big{[}D_{x}\sigma_{t_{h}}(X^{h},\L(X^{h}))\cdot\delta X^{h}+\delta_{P}\sigma_{t_{h}}(X^{h},\L(X^{h}),\tilde{X}^{h})\big{]}\mathrm{d}B_{t}$
$+\big{[}D_{x}b_{t_{h}}(X^{h},\L(X^{h}))\cdot\delta X^{h}+\delta_{P}b_{t_{h}}(X^{h},\L(X^{h}),\tilde{X}^{h})\big{]}\mathrm{d}t$
$+\mathbb{E}\big{[}D_{y}\delta_{P}b_{t_{h}}(X^{h},\L(X^{h}),X^{h})\cdot\delta X^{h}\big{|}(\tilde{B}_{s})_{0\leq s\leq t_{h}}\big{]}\mathrm{d}t$
$+\mathbb{E}\big{[}D_{y}\delta_{P}\sigma_{t_{h}}(X^{h},\L(X^{h}),X^{h})\cdot\delta X^{h}\big{|}(\tilde{B}_{s})_{0\leq s\leq t_{h}}\big{]}\mathrm{d}B_{t},\qquad\delta X^{h}_{0}=0.\phantom{abcd}$
@@ B.1 Representation formulae
if $X^{h,u}$ is $(\mathcal{F}_{t})_{t\leq 0\leq T}$-adapted and the integrated version of (31) holds $\mathbb{P}$-a.s. Let $W$ denote the Wiener measure on the canonical space $\Omega=\mathcal{C}^{d}$, with initial law Ł$\L(X^{1}_{0})$, and let $P\in\mathcal{P}_{p}(\mathcal{C}^{d})$. From (6), $\Gamma_{h}(P)$ is the law of the strong solution under $W$ to
@@
$H(P|\Gamma_{h}(P))=\inf_{\Sigma}\inf_{\mathcal{L}(X^{h,u})=P}H(\L(X^{h,u}_{0})|\L(X^{1}_{0}))+\mathbb{E}\int_{0}^{T}\frac{1}{2}|u_{t}\rvert^{2}\mathrm{d}t.$
@@
$H(R|W)=H(R_{0}|\L(X^{1}_{0}))+\inf_{\Sigma}\inf_{\mathcal{L}(Z^{u})=R}\mathbb{E}\int_{0}^{T}\frac{1}{2}|u_{t}\rvert^{2}\mathrm{d}t,$ (32)
where we minimise over processes $Z^{u}=(Z^{u}_{t})_{0\leq t\leq T}$ satisfying Ł$\L(Z^{u})=R$, and $\mathbb{P}$-a.s.
@@
$H(P|\Gamma_{h}(P))=\inf_{\Sigma}\inf_{P=Y^{h,P}_{\#}R}\inf_{\mathcal{L}(Z^{u})=R}H(R_{0}|\L(X^{1}_{0}))+\inf_{\mathcal{L}(Z^{u})=R}\mathbb{E}\int_{0}^{T}\frac{1}{2}|u_{t}\rvert^{2}\mathrm{d}t.$
@@
$\mathrm{d}X^{k}_{t}=b_{t_{h_{k}}}(X^{k},\L(X^{k}))\mathrm{d}t+\sigma_{t_{h_{k}}}(X^{k},\L(X^{k}))u^{k}_{t}\mathrm{d}t+\sigma_{t_{h_{k}}}(X^{k}_{t},\L(X^{k}))\mathrm{d}B^{k}_{t},$
@@
$\sup_{k\geq 1}H(\L(X^{k}_{0})|\L(X^{1}_{0}))+\mathbb{E}\int_{0}^{T}\frac{1}{2}\lvert u^{k}_{t}|^{2}\mathrm{d}t<+\infty.$
@@
$\sup_{0\leq s\leq T}|X^{k}_{s}\rvert^{p}\leq C\bigg{[}1+|X^{k}_{0}|^{p}+\int_{0}^{T}|u^{k}_{s}|^{p}\mathrm{d}s+\sup_{0\leq t\leq T}\bigg{|}\int_{0}^{t}\sigma_{s_{h_{k}}}(X^{k},\L(X^{k}))\mathrm{d}B^{k}_{s}\bigg{|}^{p}\bigg{]},$ (33)
@@
$\mathbb{E}\,\sup_{0\leq t\leq T}\bigg{|}\int_{0}^{t}\sigma_{s_{h_{k}}}(X^{k},\L(X^{k}))\mathrm{d}B^{k}_{s}\bigg{|}^{p}\leq\mathbb{E}\bigg{[}\bigg{(}\int_{0}^{T}|\sigma_{s_{h_{k}}}(X^{k},\L(X^{k}))|^{2}\mathrm{d}s\bigg{)}^{p/2}\bigg{]}\leq T^{p/2}M_{\sigma}.$ (34)
@@
$X^{k}_{t}-X^{k}_{s}=\int_{s}^{t}b_{r_{h_{k}}}(X^{k},\L(X^{k}))\mathrm{d}r+\int_{s}^{t}\sigma_{r_{h_{k}}}(X^{k},\L(X^{k}))u^{k}_{r}\mathrm{d}r+\int_{s}^{t}\sigma_{r_{h_{k}}}(X^{k},\L(X^{k}))\mathrm{d}B^{k}_{r}.$
@@
$\mathbb{E}\bigg{|}\int_{s}^{t}\sigma_{r_{h_{k}}}(X^{k},\L(X^{k}))u^{k}_{r}\mathrm{d}r\bigg{|}\leq(t-s)^{1/2}M_{\sigma}\mathbb{E}\bigg{[}\int_{0}^{T}|u^{k}_{r}\rvert^{2}\mathrm{d}r\bigg{]}.$
@@
$H(\L(X^{k}_{0})|\L(X^{1}_{0}))=\sup_{\begin{subarray}{c}\phi\text{ measurable}\\ \mathbb{E}[\phi(X^{1}_{0})]<+\infty\end{subarray}}\mathbb{E}[\phi(X^{k}_{0})]-\log\mathbb{E}[e^{\phi(X^{1}_{0})}],$
Applying this to $\phi(x)=\alpha|x|^{p}\mathbbm{1}_{\lvert\phi\rvert\geq M}$ for every $\alpha,M>0$, uniform integrability for $(|X^{k}_{0}|^{p})_{k\geq 1}$ follows from (A.3) and the bound on $(H(\L(X^{k}_{0})|\L(X^{1}_{0})))_{k\geq 1}$. For the uniform integrability of
$\int_{0}^{T}|u^{k}_{s}|^{p}\mathrm{d}s\quad\text{and}\quad\sup_{0\leq t\leq T}\bigg{|}\int_{0}^{t}\sigma_{s_{h_{k}}}(X^{k},\L(X^{k}))\mathrm{d}B^{k}_{s}\bigg{|}^{p},$
@@
$\int_{0}^{T}|u^{k}_{s}|^{2}\mathrm{d}s\quad\text{and}\quad\sup_{0\leq t\leq T}\bigg{|}\int_{0}^{t}\sigma_{s_{h_{k}}}(X^{k},\L(X^{k}))\mathrm{d}B^{k}_{s}\bigg{|}^{2}$
@@
$\mathrm{d}X^{h,k}_{t}=b_{t_{h}}(X^{h,k},\L(X^{h,k}))\mathrm{d}t+\sigma_{t_{h}}(X^{h,k},\L(X^{h,k}))u^{k}_{t}\mathrm{d}t+\sigma_{t_{h}}(X^{h,k}_{t},\L(X^{h,k}))\mathrm{d}B^{k}_{t},$
@@
$\sup_{k\geq 1}H(\L(X^{k}_{0})|\L(X^{1}_{0}))+\mathbb{E}\int_{0}^{T}\frac{1}{2}\lvert u^{k}_{t}|^{2}\mathrm{d}t<+\infty.$
@@ B.2 Gamma-convergence
$\inf_{P\in\mathcal{P}_{p}(\mathcal{C}^{d})}H(P|\Gamma_{h}(P))+F(P)=\inf_{\Sigma,u,X^{u}_{0}}F(\L(X^{h,u}))+H(\L(X^{u}_{0})|\L(X^{1}_{0}))+\mathbb{E}\int_{0}^{T}\frac{1}{2}|u_{t}\rvert^{2}\mathrm{d}t.$
@@
$\mathrm{d}X^{h,M}_{t}=b_{t_{h}}(X^{h,M},\L(X^{h,M}))\mathrm{d}t+\sigma_{t_{h}}(X^{h,M},\L(X^{h,M}))u^{M}_{t}\mathrm{d}t+\sigma_{t_{h}}(X^{h,M},\L(X^{h,M}))\mathrm{d}B^{P}_{t},$
with $X^{h,M}_{0}=X^{0}$, has a pathwise-unique strong solution under $P$. Lemma B.2 shows that (Ł$(\L_{P}(X^{h,M}))_{h\geq 0}$ is pre-compact in $\mathcal{P}_{p}(\mathcal{C}^{d})$. From any sub-sequence of (Ł$(\L_{P}(X^{h,M}))_{h\geq 0}$, we can thus extract a sub-sequence that converges towards some $P^{M}$ in $\mathcal{P}_{p}(\mathcal{C}^{d})$. Up to reindexing, let us assume that Ł$\L_{P}(X^{h,M})\rightarrow P^{M}$ as $h\rightarrow 0$. Using the Girsanov transform, the SDE
@@
$\mathrm{d}X^{h,M}_{t}=b_{t_{h}}(X^{h,M},\L_{P}(X^{h,M}))\mathrm{d}t+\sigma_{t_{h}}(X^{h,M},\L_{P}(X^{h,M}))\mathrm{d}B^{Q}_{t},\quad X^{h,M}_{0}=X^{0},$
@@
these equations holding $Q$-.a.s. As previously, a coupling argument yields the stability result $\mathbb{E}_{Q}[\sup_{0\leq t\leq T}|X^{h,M}_{t}-X^{M}_{t}|^{p}]\rightarrow 0$ as $h\rightarrow 0$. By (35) and dominated convergence, we deduce that Ł$\L_{P}(X^{h,M})$ weakly converges towards Ł$\L_{P}(X^{M})$, proving that $P^{M}=\L_{P}(X^{M})$. Thus, $X^{M}$ is the strong solution of a McKean-Vlasov equation under $P$.
@@
$(W_{p}(\L(X^{0}),\L(X^{M})))^{2}\leq C\mathbb{E}\big{[}\sup_{0\leq t\leq T}|X^{h,M}_{t}-X^{M}_{t}|^{2}\big{]},$
and we can now use coupling arguments in $L^{2}$. At this stage, we use the definition of $u^{M}$ and the proof of [fischer2014form, Proposition C.1] to obtain that $\mathbb{E}_{P}[\sup_{0\leq t\leq T}|X^{0}_{t\wedge\tau_{M}}-X^{M}_{t\wedge\tau_{M}}|^{2}]=0$, where $\tau_{M}:=\inf\{t\in[0,T],\,\mathcal{E}_{t}(u)\geq M\}$. From Lemma B.2, (Ł$(\L_{P}(X^{M}))_{M>0}$ is pre-compact in $\mathcal{P}_{p}(\mathcal{C}^{d})$. Sending $M\rightarrow+\infty$, we eventually deduce that $P^{M}=\L_{P}(X^{M})$ converges towards $P=\L_{P}(X^{0})$.
@@
¯$\overline{\mathcal{F}}_{h}\leq F(\L(X^{h,M}))+\mathbb{E}_{P}\int_{0}^{T}|u^{M}_{t}|^{2}\mathrm{d}t.$

### v2

@@ 4.3 Uniform fluctuation estimates
$\delta^{h,i}_{t}=\int_{0}^{t}[b_{s_{h}}(X^{h,i},\pi(\vec{X}^{h}))-b_{s_{h}}(\overline{X}^{h,i},\mathcal{L}(\overline{X}^{h,i}))]\mathrm{d}s\\ +\int_{0}^{t}[\sigma_{s_{h}}(X^{h,i},\pi(\vec{X}^{h}))-\sigma_{s_{h}}(\overline{X}^{h,i},\mathcal{L}(\overline{X}^{h,i}))]\mathrm{d}B^{i}_{s}.$
@@
$\mathbb{E}\big{|}b_{s_{h}}(X^{h,i},\pi(\vec{X}^{h}))-b_{s_{h}}(\overline{X}^{h,i},\mathcal{L}(\overline{X}^{h,i}))\big{|}^{2}\leq C\mathbb{E}\big{[}\sup_{0\leq r\leq s_{h}}|X^{h,i}_{r}-\overline{X}^{h,i}_{r}|^{2}\big{]}+CN^{-1}.$
@@
$\mathbb{E}\sup_{0\leq s\leq t}\bigg{|}\int_{0}^{s}[\sigma_{r_{h}}(X^{h,i},\pi(\vec{X}^{h}))-\sigma_{r_{h}}(\overline{X}^{h,i},\mathcal{L}(\overline{X}^{h,i}))]\mathrm{d}B^{i}_{r}\bigg{|}^{2}\\ \leq C\int_{0}^{t}|\sigma_{s_{h}}(X^{h,i},\pi(\vec{X}^{h}))-\sigma_{s_{h}}(\overline{X}^{h,i},\mathcal{L}(\overline{X}^{h,i}))|^{2}\mathrm{d}s,$
@@
$b_{s}^{h,i}:=b_{s_{h}}(X^{h,i},\pi(\vec{X}^{h}))-b_{s_{h}}(\overline{X}^{h,i},\mathcal{L}(\overline{X}^{h,i}))-[b_{s}(X^{0,i},\pi(\vec{X}^{\infty}))-b_{s}(\overline{X}^{\infty,i},\mathcal{L}(\overline{X}^{0,i}))],$
$\sigma_{s}^{h,i}:=\sigma_{s_{h}}(X^{h,i},\pi(\vec{X}^{h}))-\sigma_{s_{h}}(\overline{X}^{h,i},\mathcal{L}(\overline{X}^{h,i}))-[\sigma_{s}(X^{0,i},\pi(\vec{X}^{\infty}))-\sigma_{s}(\overline{X}^{\infty,i},\mathcal{L}(\overline{X}^{0,i}))].$
@@
$+b_{s_{h}}(\overline{X}^{h,i},\pi(\vec{\overline{X}}^{h}))-b_{s_{h}}(\overline{X}^{h,i},\mathcal{L}(\overline{X}^{h,i}))-\big{[}b_{s}(\overline{X}^{h,i},\pi(\vec{\overline{X}}^{h}))-b_{s}(\overline{X}^{h,i},\mathcal{L}(\overline{X}^{h,i}))\big{]}$
@@
$+b_{s}(\overline{X}^{h,i},\pi(\vec{\overline{X}}^{h}))-b_{s}(\overline{X}^{h,i},\mathcal{L}(\overline{X}^{h,i}))-\big{[}b_{s}(\overline{X}^{i,0}_{s},\pi(\vec{\overline{X}}^{0}_{s}))-b_{s}(\overline{X}^{i,0},\mathcal{L}(\overline{X}^{i,0}))\big{]}.$
@@
$\max_{0\leq r\leq 1}\mathbb{E}\big{[}\big{|}\partial_{s}b_{(1-r)s_{h}+rs}(\overline{X}^{h,i},\pi(\vec{\overline{X}}^{h}))-\partial_{s}b_{(1-r)s_{h}+rs}(\overline{X}^{h,i},\mathcal{L}(\overline{X}^{h,i}))\big{|}\big{]}\leq CN^{-1/2}.$
@@
$\mathrm{d}\delta X^{h}_{t_{h}}$ $=\big{[}D_{x}\sigma_{t_{h}}(X^{h},\mathcal{L}(X^{h}))\cdot\delta X^{h}+\delta_{P}\sigma_{t_{h}}(X^{h},\mathcal{L}(X^{h}),\tilde{X}^{h})\big{]}\mathrm{d}B_{t}$
$+\big{[}D_{x}b_{t_{h}}(X^{h},\mathcal{L}(X^{h}))\cdot\delta X^{h}+\delta_{P}b_{t_{h}}(X^{h},\mathcal{L}(X^{h}),\tilde{X}^{h})\big{]}\mathrm{d}t$
$+\mathbb{E}\big{[}D_{y}\delta_{P}b_{t_{h}}(X^{h},\mathcal{L}(X^{h}),X^{h})\cdot\delta X^{h}\big{|}(\tilde{B}_{s})_{0\leq s\leq t_{h}}\big{]}\mathrm{d}t$
$+\mathbb{E}\big{[}D_{y}\delta_{P}\sigma_{t_{h}}(X^{h},\mathcal{L}(X^{h}),X^{h})\cdot\delta X^{h}\big{|}(\tilde{B}_{s})_{0\leq s\leq t_{h}}\big{]}\mathrm{d}B_{t},\qquad\delta X^{h}_{0}=0.\phantom{abcd}$
@@ B.1 Representation formulae
if $X^{h,u}$ is $(\mathcal{F}_{t})_{t\leq 0\leq T}$-adapted and the integrated version of (31) holds $\mathbb{P}$-a.s. Let $W$ denote the Wiener measure on the canonical space $\Omega=\mathcal{C}^{d}$, with initial law $\mathcal{L}(X^{1}_{0})$, and let $P\in\mathcal{P}_{p}(\mathcal{C}^{d})$. From (6), $\Gamma_{h}(P)$ is the law of the strong solution under $W$ to
@@
$H(P|\Gamma_{h}(P))=\inf_{\Sigma}\inf_{\mathcal{L}(X^{h,u})=P}H(\mathcal{L}(X^{h,u}_{0})|\mathcal{L}(X^{1}_{0}))+\mathbb{E}\int_{0}^{T}\frac{1}{2}|u_{t}\rvert^{2}\mathrm{d}t.$
@@
$H(R|W)=H(R_{0}|\mathcal{L}(X^{1}_{0}))+\inf_{\Sigma}\inf_{\mathcal{L}(Z^{u})=R}\mathbb{E}\int_{0}^{T}\frac{1}{2}|u_{t}\rvert^{2}\mathrm{d}t,$ (32)
where we minimise over processes $Z^{u}=(Z^{u}_{t})_{0\leq t\leq T}$ satisfying $\mathcal{L}(Z^{u})=R$, and $\mathbb{P}$-a.s.
@@
$H(P|\Gamma_{h}(P))=\inf_{\Sigma}\inf_{P=Y^{h,P}_{\#}R}\inf_{\mathcal{L}(Z^{u})=R}H(R_{0}|\mathcal{L}(X^{1}_{0}))+\inf_{\mathcal{L}(Z^{u})=R}\mathbb{E}\int_{0}^{T}\frac{1}{2}|u_{t}\rvert^{2}\mathrm{d}t.$
@@
$\mathrm{d}X^{k}_{t}=b_{t_{h_{k}}}(X^{k},\mathcal{L}(X^{k}))\mathrm{d}t+\sigma_{t_{h_{k}}}(X^{k},\mathcal{L}(X^{k}))u^{k}_{t}\mathrm{d}t+\sigma_{t_{h_{k}}}(X^{k}_{t},\mathcal{L}(X^{k}))\mathrm{d}B^{k}_{t},$
@@
$\sup_{k\geq 1}H(\mathcal{L}(X^{k}_{0})|\mathcal{L}(X^{1}_{0}))+\mathbb{E}\int_{0}^{T}\frac{1}{2}\lvert u^{k}_{t}|^{2}\mathrm{d}t<+\infty.$
@@
$\sup_{0\leq s\leq T}|X^{k}_{s}\rvert^{p}\leq C\bigg{[}1+|X^{k}_{0}|^{p}+\int_{0}^{T}|u^{k}_{s}|^{p}\mathrm{d}s+\sup_{0\leq t\leq T}\bigg{|}\int_{0}^{t}\sigma_{s_{h_{k}}}(X^{k},\mathcal{L}(X^{k}))\mathrm{d}B^{k}_{s}\bigg{|}^{p}\bigg{]},$ (33)
@@
$\mathbb{E}\,\sup_{0\leq t\leq T}\bigg{|}\int_{0}^{t}\sigma_{s_{h_{k}}}(X^{k},\mathcal{L}(X^{k}))\mathrm{d}B^{k}_{s}\bigg{|}^{p}\leq\mathbb{E}\bigg{[}\bigg{(}\int_{0}^{T}|\sigma_{s_{h_{k}}}(X^{k},\mathcal{L}(X^{k}))|^{2}\mathrm{d}s\bigg{)}^{p/2}\bigg{]}\leq T^{p/2}M_{\sigma}.$ (34)
@@
$X^{k}_{t}-X^{k}_{s}=\int_{s}^{t}b_{r_{h_{k}}}(X^{k},\mathcal{L}(X^{k}))\mathrm{d}r+\int_{s}^{t}\sigma_{r_{h_{k}}}(X^{k},\mathcal{L}(X^{k}))u^{k}_{r}\mathrm{d}r+\int_{s}^{t}\sigma_{r_{h_{k}}}(X^{k},\mathcal{L}(X^{k}))\mathrm{d}B^{k}_{r}.$
@@
$\mathbb{E}\bigg{|}\int_{s}^{t}\sigma_{r_{h_{k}}}(X^{k},\mathcal{L}(X^{k}))u^{k}_{r}\mathrm{d}r\bigg{|}\leq(t-s)^{1/2}M_{\sigma}\mathbb{E}\bigg{[}\int_{0}^{T}|u^{k}_{r}\rvert^{2}\mathrm{d}r\bigg{]}.$
@@
$H(\mathcal{L}(X^{k}_{0})|\mathcal{L}(X^{1}_{0}))=\sup_{\begin{subarray}{c}\phi\text{ measurable}\\ \mathbb{E}[\phi(X^{1}_{0})]<+\infty\end{subarray}}\mathbb{E}[\phi(X^{k}_{0})]-\log\mathbb{E}[e^{\phi(X^{1}_{0})}],$
Applying this to $\phi(x)=\alpha|x|^{p}\mathbbm{1}_{\lvert\phi\rvert\geq M}$ for every $\alpha,M>0$, uniform integrability for $(|X^{k}_{0}|^{p})_{k\geq 1}$ follows from (A.3) and the bound on $(H(\mathcal{L}(X^{k}_{0})|\mathcal{L}(X^{1}_{0})))_{k\geq 1}$. For the uniform integrability of
$\int_{0}^{T}|u^{k}_{s}|^{p}\mathrm{d}s\quad\text{and}\quad\sup_{0\leq t\leq T}\bigg{|}\int_{0}^{t}\sigma_{s_{h_{k}}}(X^{k},\mathcal{L}(X^{k}))\mathrm{d}B^{k}_{s}\bigg{|}^{p},$
@@
$\int_{0}^{T}|u^{k}_{s}|^{2}\mathrm{d}s\quad\text{and}\quad\sup_{0\leq t\leq T}\bigg{|}\int_{0}^{t}\sigma_{s_{h_{k}}}(X^{k},\mathcal{L}(X^{k}))\mathrm{d}B^{k}_{s}\bigg{|}^{2}$
@@
$\mathrm{d}X^{h,k}_{t}=b_{t_{h}}(X^{h,k},\mathcal{L}(X^{h,k}))\mathrm{d}t+\sigma_{t_{h}}(X^{h,k},\mathcal{L}(X^{h,k}))u^{k}_{t}\mathrm{d}t+\sigma_{t_{h}}(X^{h,k}_{t},\mathcal{L}(X^{h,k}))\mathrm{d}B^{k}_{t},$
@@
$\sup_{k\geq 1}H(\mathcal{L}(X^{k}_{0})|\mathcal{L}(X^{1}_{0}))+\mathbb{E}\int_{0}^{T}\frac{1}{2}\lvert u^{k}_{t}|^{2}\mathrm{d}t<+\infty.$
@@ B.2 Gamma-convergence
$\inf_{P\in\mathcal{P}_{p}(\mathcal{C}^{d})}H(P|\Gamma_{h}(P))+F(P)=\inf_{\Sigma,u,X^{u}_{0}}F(\mathcal{L}(X^{h,u}))+H(\mathcal{L}(X^{u}_{0})|\mathcal{L}(X^{1}_{0}))+\mathbb{E}\int_{0}^{T}\frac{1}{2}|u_{t}\rvert^{2}\mathrm{d}t.$
@@
$\mathrm{d}X^{h,M}_{t}=b_{t_{h}}(X^{h,M},\mathcal{L}(X^{h,M}))\mathrm{d}t+\sigma_{t_{h}}(X^{h,M},\mathcal{L}(X^{h,M}))u^{M}_{t}\mathrm{d}t+\sigma_{t_{h}}(X^{h,M},\mathcal{L}(X^{h,M}))\mathrm{d}B^{P}_{t},$
with $X^{h,M}_{0}=X^{0}$, has a pathwise-unique strong solution under $P$. Lemma B.2 shows that $(\mathcal{L}_{P}(X^{h,M}))_{h\geq 0}$ is pre-compact in $\mathcal{P}_{p}(\mathcal{C}^{d})$. From any sub-sequence of $(\mathcal{L}_{P}(X^{h,M}))_{h\geq 0}$, we can thus extract a sub-sequence that converges towards some $P^{M}$ in $\mathcal{P}_{p}(\mathcal{C}^{d})$. Up to reindexing, let us assume that $\mathcal{L}_{P}(X^{h,M})\rightarrow P^{M}$ as $h\rightarrow 0$. Using the Girsanov transform, the SDE
@@
$\mathrm{d}X^{h,M}_{t}=b_{t_{h}}(X^{h,M},\mathcal{L}_{P}(X^{h,M}))\mathrm{d}t+\sigma_{t_{h}}(X^{h,M},\mathcal{L}_{P}(X^{h,M}))\mathrm{d}B^{Q}_{t},\quad X^{h,M}_{0}=X^{0},$
@@
these equations holding $Q$-.a.s. As previously, a coupling argument yields the stability result $\mathbb{E}_{Q}[\sup_{0\leq t\leq T}|X^{h,M}_{t}-X^{M}_{t}|^{p}]\rightarrow 0$ as $h\rightarrow 0$. By (35) and dominated convergence, we deduce that $\mathcal{L}_{P}(X^{h,M})$ weakly converges towards $\mathcal{L}_{P}(X^{M})$, proving that $P^{M}=\mathcal{L}_{P}(X^{M})$. Thus, $X^{M}$ is the strong solution of a McKean-Vlasov equation under $P$.
@@
$(W_{p}(\mathcal{L}(X^{0}),\mathcal{L}(X^{M})))^{2}\leq C\mathbb{E}\big{[}\sup_{0\leq t\leq T}|X^{h,M}_{t}-X^{M}_{t}|^{2}\big{]},$
and we can now use coupling arguments in $L^{2}$. At this stage, we use the definition of $u^{M}$ and the proof of [fischer2014form, Proposition C.1] to obtain that $\mathbb{E}_{P}[\sup_{0\leq t\leq T}|X^{0}_{t\wedge\tau_{M}}-X^{M}_{t\wedge\tau_{M}}|^{2}]=0$, where $\tau_{M}:=\inf\{t\in[0,T],\,\mathcal{E}_{t}(u)\geq M\}$. From Lemma B.2, $(\mathcal{L}_{P}(X^{M}))_{M>0}$ is pre-compact in $\mathcal{P}_{p}(\mathcal{C}^{d})$. Sending $M\rightarrow+\infty$, we eventually deduce that $P^{M}=\mathcal{L}_{P}(X^{M})$ converges towards $P=\mathcal{L}_{P}(X^{0})$.
@@
¯$\overline{\mathcal{F}}_{h}\leq F(\mathcal{L}(X^{h,M}))+\mathbb{E}_{P}\int_{0}^{T}|u^{M}_{t}|^{2}\mathrm{d}t.$